$SPAD/src/input equation2.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool equation2.output
)set message test on
)set message auto off
)clear all
--S 1 of 27
solve([3*x**3 + y + 1,y - 1],[x,y])
--R
--R
--R            3
--R   (1)  [[3x  + 2= 0,y= 1]]
--R                    Type: List(List(Equation(Fraction(Polynomial(Integer)))))
--E 1
--S 2 of 27
solve([x**3 + x - y**2 + 4,x*y + 2],[x,y],"sym")
--R
--R   There are 6 exposed and 1 unexposed library operations named solve
--R      having 3 argument(s) but none was determined to be applicable.
--R      Use HyperDoc Browse, or issue
--R                              )display op solve
--R      to learn more about the available operations. Perhaps
--R      package-calling the operation or using coercions on the arguments
--R      will allow you to apply the operation.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named solve
--R      with argument type(s)
--R                          List(Polynomial(Integer))
--R                      List(OrderedVariableList([x,y]))
--R                                   String
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 2
--S 3 of 27
solve([x = y**2-19,y = z**2+x+3,z = 3*x],[x,y,z])
--R

--R
--R                    2
--R             z    3z  + z + 9   4     3      2
--R   (2)  [[x= -,y= -----------,9z  + 6z  + 55z  + 15z - 90= 0]]
--R             3         3
--R                    Type: List(List(Equation(Fraction(Polynomial(Integer)))))
--E 3
--S 4 of 27
solve([3*x + 2*y - z,x - 1/2*y + 1/3*z,4/5*x - 2/3*y - z])
--R
--R
--R   (3)  [[z= 0,y= 0,x= 0]]
--R                    Type: List(List(Equation(Fraction(Polynomial(Integer)))))
--E 4
--S 5 of 27
solve([x**2*y - 1,x*y**2 - 2],[x,y],.01)
--R
--R   There are 6 exposed and 1 unexposed library operations named solve
--R      having 3 argument(s) but none was determined to be applicable.
--R      Use HyperDoc Browse, or issue
--R                              )display op solve
--R      to learn more about the available operations. Perhaps
--R      package-calling the operation or using coercions on the arguments
--R      will allow you to apply the operation.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named solve
--R      with argument type(s)
--R                          List(Polynomial(Integer))
--R                      List(OrderedVariableList([x,y]))
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 5
--S 6 of 27
solve([x**2/a = 1,a**2 - a*x = 0],[x,a],.001)
--R
--R   There are 6 exposed and 1 unexposed library operations named solve
--R      having 3 argument(s) but none was determined to be applicable.
--R      Use HyperDoc Browse, or issue
--R                              )display op solve
--R      to learn more about the available operations. Perhaps
--R      package-calling the operation or using coercions on the arguments
--R      will allow you to apply the operation.
--R
--RDaly Bug

--R   Cannot find a definition or applicable library operation named solve
--R      with argument type(s)
--R                List(Equation(Fraction(Polynomial(Integer))))
--R                      List(OrderedVariableList([x,a]))
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 6
--S 7 of 27
solve([x**2/a + a + y**3 - 1,a*y + a + 1],[x,y])
--R
--R
--R           2 2    4     3     2                - a - 1
--R   (4)  [[a x  + a  - 2a  - 3a  - 3a - 1= 0,y= -------]]
--R                                                  a
--R                    Type: List(List(Equation(Fraction(Polynomial(Integer)))))
--E 7
)clear all
--S 8 of 27
solve(x**3 + 1 = 0,x)
--R
--R
--R                 2
--R   (1)  [x= - 1,x  - x + 1= 0]
--R                          Type: List(Equation(Fraction(Polynomial(Integer))))
--E 8
--S 9 of 27
solve(x**3*y + x*y + 1,x,"sym")
--R
--R   There are 6 exposed and 1 unexposed library operations named solve
--R      having 3 argument(s) but none was determined to be applicable.
--R      Use HyperDoc Browse, or issue
--R                              )display op solve
--R      to learn more about the available operations. Perhaps
--R      package-calling the operation or using coercions on the arguments
--R      will allow you to apply the operation.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named solve
--R      with argument type(s)
--R                             Polynomial(Integer)
--R                                 Variable(x)
--R                                   String
--R
--R      Perhaps you should use "@" to indicate the required return type,

--R      or "$" to specify which version of the function you need.
--E 9
--S 10 of 27
solve(3*x + 1/4*y = 1,x)
--R
--R
--R            - y + 4
--R   (2)  [x= -------]
--R               12
--R                          Type: List(Equation(Fraction(Polynomial(Integer))))
--E 10
--S 11 of 27
solve(x**4 - 10*x**3 + 35*x**2 - 50*x + 25,x,1/1000)
--R
--R   There are 6 exposed and 1 unexposed library operations named solve
--R      having 3 argument(s) but none was determined to be applicable.
--R      Use HyperDoc Browse, or issue
--R                              )display op solve
--R      to learn more about the available operations. Perhaps
--R      package-calling the operation or using coercions on the arguments
--R      will allow you to apply the operation.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named solve
--R      with argument type(s)
--R                             Polynomial(Integer)
--R                                 Variable(x)
--R                              Fraction(Integer)
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 11
--S 12 of 27
solve(x**4 - 10*x**3 + 35*x**2 - 50*x + 25,x,"sym")
--R
--R   There are 6 exposed and 1 unexposed library operations named solve
--R      having 3 argument(s) but none was determined to be applicable.
--R      Use HyperDoc Browse, or issue
--R                              )display op solve
--R      to learn more about the available operations. Perhaps
--R      package-calling the operation or using coercions on the arguments
--R      will allow you to apply the operation.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named solve
--R      with argument type(s)
--R                             Polynomial(Integer)

--R                                 Variable(x)
--R                                   String
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 12
--S 13 of 27
solve(x**3 - sqrt(2))
--R
--R
--R          3    +-+
--R   (3)  [x  - \|2 = 0]
--R                  Type: List(Equation(Fraction(Polynomial(AlgebraicNumber))))
--E 13
--S 14 of 27
solve(x**3/a + x/a + 1,x)
--R
--R
--R          3
--R   (4)  [x  + x + a= 0]
--R                          Type: List(Equation(Fraction(Polynomial(Integer))))
--E 14
)clear all
--S 15 of 27
solve(1/x**3 + 1/x**2 + 1/x = 0,x,"sym")
--R
--R   There are 6 exposed and 1 unexposed library operations named solve
--R      having 3 argument(s) but none was determined to be applicable.
--R      Use HyperDoc Browse, or issue
--R                              )display op solve
--R      to learn more about the available operations. Perhaps
--R      package-calling the operation or using coercions on the arguments
--R      will allow you to apply the operation.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named solve
--R      with argument type(s)
--R                   Equation(Fraction(Polynomial(Integer)))
--R                                 Variable(x)
--R                                   String
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 15
--S 16 of 27

solve(x**3 + 1 = 0,x)
--R
--R
--R                 2
--R   (1)  [x= - 1,x  - x + 1= 0]
--R                          Type: List(Equation(Fraction(Polynomial(Integer))))
--E 16
--S 17 of 27
solve(x**3*y + x*y + 1,x,"sym")
--R
--R   There are 6 exposed and 1 unexposed library operations named solve
--R      having 3 argument(s) but none was determined to be applicable.
--R      Use HyperDoc Browse, or issue
--R                              )display op solve
--R      to learn more about the available operations. Perhaps
--R      package-calling the operation or using coercions on the arguments
--R      will allow you to apply the operation.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named solve
--R      with argument type(s)
--R                             Polynomial(Integer)
--R                                 Variable(x)
--R                                   String
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 17
--S 18 of 27
solve(3*x + 1/4*y = 1,x)
--R
--R
--R            - y + 4
--R   (2)  [x= -------]
--R               12
--R                          Type: List(Equation(Fraction(Polynomial(Integer))))
--E 18
--S 19 of 27
solve(x**4 - 10*x**3 + 35*x**2 - 50*x + 25,x,1/1000)
--R
--R   There are 6 exposed and 1 unexposed library operations named solve
--R      having 3 argument(s) but none was determined to be applicable.
--R      Use HyperDoc Browse, or issue
--R                              )display op solve
--R      to learn more about the available operations. Perhaps
--R      package-calling the operation or using coercions on the arguments
--R      will allow you to apply the operation.

--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named solve
--R      with argument type(s)
--R                             Polynomial(Integer)
--R                                 Variable(x)
--R                              Fraction(Integer)
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 19
--S 20 of 27
solve(x**4 - 10*x**3 + 35*x**2 - 50*x + 25,x,"sym")
--R
--R   There are 6 exposed and 1 unexposed library operations named solve
--R      having 3 argument(s) but none was determined to be applicable.
--R      Use HyperDoc Browse, or issue
--R                              )display op solve
--R      to learn more about the available operations. Perhaps
--R      package-calling the operation or using coercions on the arguments
--R      will allow you to apply the operation.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named solve
--R      with argument type(s)
--R                             Polynomial(Integer)
--R                                 Variable(x)
--R                                   String
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 20
--S 21 of 27
solve(x**3 - sqrt(2))
--R
--R
--R          3    +-+
--R   (3)  [x  - \|2 = 0]
--R                  Type: List(Equation(Fraction(Polynomial(AlgebraicNumber))))
--E 21
--S 22 of 27
solve(x**3/a + x/a + 1,x)
--R
--R
--R          3
--R   (4)  [x  + x + a= 0]
--R                          Type: List(Equation(Fraction(Polynomial(Integer))))

--E 22
--S 23 of 27
solve(1/x**3 + 1/x**2 + 1/x = 0,x,"sym")
--R
--R   There are 6 exposed and 1 unexposed library operations named solve
--R      having 3 argument(s) but none was determined to be applicable.
--R      Use HyperDoc Browse, or issue
--R                              )display op solve
--R      to learn more about the available operations. Perhaps
--R      package-calling the operation or using coercions on the arguments
--R      will allow you to apply the operation.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named solve
--R      with argument type(s)
--R                   Equation(Fraction(Polynomial(Integer)))
--R                                 Variable(x)
--R                                   String
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 23
)clear all
--S 24 of 27
solve([[1,1,1],[3,-2,1],[1,2,2]],[8,0,17])
--R
--R
--R   (1)  [particular= [- 1,2,7],basis= [[0,0,0]]]
--IType: Record(particular: Union(Vector(Fraction(Integer)),"failed"),...
--E 24
--S 25 of 27
solve([[1,2,3],[2,3,4],[3,4,5]],[2,2,2])
--R
--R
--R   (2)  [particular= [- 2,2,0],basis= [[1,- 2,1]]]
--IType: Record(particular: Union(Vector(Fraction(Integer)),"failed"),...
--E 25
--S 26 of 27
solve([[1,2,3],[2,3,4],[3,4,5]],[2,3,2])
--R
--R
--R   (3)  [particular= "failed",basis= [[1,- 2,1]]]
--IType: Record(particular: Union(Vector(Fraction(Integer)),"failed"),...
--E 26

--S 27 of 27
solve([[1,2,3],[2,3,4],[3,4,5]])
--R
--R
--R   (4)  solve
--R             [1,2,3],[2,3,4],[3,4,5]
--R                                                                 Type: Symbol
--E 27
)spool
)lisp (bye)

———-

References^{[1] nothing}

12